C. Let p and q be proposition

p: 4 is a rational number

q: √3 is an irrational number


Express each of these proposition as an English sentences:


  1. ~p
  2. p∨ q
  3. p ∧ q
  4. p → q
  5. p ↔ q
1
Expert's answer
2022-03-29T02:02:18-0400

1: 4 is not a rational number

2: 4 is a rational number or "\sqrt{3}" is an irrational number

3: 4 is a rational number and "\sqrt{3}" is an irrational number

4: if 4 is a rational number then "\sqrt{3}" is an irrational number

5: 4 is a rational number if and only if "\sqrt{3}" is an irrational number